# Efficient Unsupervised Domain Adaptation Regression for Spatial-Temporal Air Quality Sensor Fusion

### v1

@@ 3.1 TikUDA
The HS values between the subspace spans of the source and target features are assigned to $\mathbf{m}=[\mathsf{HS}(\phi^{S\leftrightarrow T}_{(1)}),\ldots,\mathsf{HS}(\phi^{S\leftrightarrow T}_{(p)})]$. Thus, the loss function to align the bases from the inverse Tikhonov matrix can be expressed as:
@@
3.2 Spatial-Temporal Feature Extractor
@@
Figure 3: An overview of the proposed approach for domain adaptive regression problems on spatial-temporal graph data
Temporal Module of Feature Extractor : Temporal dependencies, also known as temporal patterns, can be effectively represented in both the time domain and the frequency domain [34]. To address these dependencies, we incorporate temporal modules before spatial modules, facilitating the modeling of complex spatial-temporal patterns. Our approach adopts recurrence-based methodologies, specifically leveraging the Gated Recurrent Unit (GRU), to capture and interpret the underlying temporal patterns inherent in the time series data by:
$\mathbf{r_{t}}$ $=\sigma(\mathbf{W}_{\text{r}}\mathbf{x_{t}}+\mathbf{U}_{\text{r}}\mathbf{h_{t-1}}+\mathbf{b}_{\text{r}})$ (8)
$\mathbf{z_{t}}$ $=\sigma(\mathbf{W}_{\text{z}}\mathbf{x_{t}}+\mathbf{U}_{\text{z}}\mathbf{h_{t-1}}+\mathbf{b}_{\text{z}})$ (9)
$\mathbf{\hat{h}_{t}}$ $=\tanh(\mathbf{W}_{\text{h}}\mathbf{x_{t}}+\mathbf{U}_{\text{h}}(\mathbf{r}_{\text{t}}\odot\mathbf{h_{t-1}})+\mathbf{b}_{\text{h}})$ (10)
$\mathbf{h_{t}}$ $=(1-\mathbf{z_{t}})\odot\mathbf{h_{t-1}}+\mathbf{z_{t}}\odot\mathbf{\hat{h}_{t}}$ (11)
In this formula, $\odot$ represents the Hadamard product operator, $\mathbf{x_{t}}$ is the input vector, $\mathbf{r_{t}}$ represents the reset gate vector, which controls how much of the previous state should be ignored, $\mathbf{z_{t}}$ represents the update gate, which determines how much of the previous state should be passed along, $\mathbf{\hat{h}_{t}}$ represents the new gate, which computes the new candidate values for the memory cell, and $\mathbf{h}$ is the output vector.
@@
where $\mathbf{a}$ represents a trainable attention vector, $\mathbf{W}$ is a trainable matrix, and ||||| | denotes the concatenation operation. For brevity, the $k$-th GAT layer can be represented compactly as:
@@
3.3 Overall Framework and Summary
@@
Next, the data undergoes a temporal processing step using GRU to identify and extract temporal patterns. This is followed by a spatial processing step employing a GAT, leveraging a fully connected graph due to the absence of a predefined graph structure. The extracted features are then inputted into a regressor to predict the quantities of pollutants.
The training objective is to minimize both the prediction error on the source domain and the alignment of features between the source and target domains. The total loss function can be expressed as:
@@
The loss term $L_{\text{src}}$ represents the supervised loss (mean squared error) on the source domain, comparing the prediction with the ground truth. The terms $L_{\text{angle}}$ and $L_{\text{scale}}$ account for the angle and scale alignment with respect to the inverse Tikhonov matrix. The hyperparameters $\gamma_{\text{angle}}$ and $\gamma_{\text{scale}}$ control the influence of angle and scale alignment on the total loss. Finally, for a detailed representation of the domain adaptation part from source to target, Figure 4 illustrates the flowchart of TikUDA. In summary, TikUDA is designed to align the inverse of the Gram matrix of features, which is perturbed by a scaled identity matrix referred to as the Tikhonov matrix. This perturbation ensures that the Gram matrix becomes full rank and invertible. A key aspect of the method is that the Tikhonov matrix is both symmetric and positive definite, which allows for the use of Cholesky decomposition—a more computationally efficient approach to matrix inversion compared to SVD, particularly for large matrices. Cholesky decomposition, specifically suited for symmetric and positive definite matrices, enhances numerical stability while maintaining the structural properties of the Tikhonov matrix. Additionally, TikUDA introduces novel similarity metrics (Haversine similarity) as regularization terms to improve subspaces’ alignment between the source and target feature domains. Moreover, TikUDA focuses on scale alignment between source and target domains by aligning the largest principal eigenvalues of the inverse Tikhonov matrices, as the largest eigenvalue demostrates the greatest influence on feature scaling. Instead of computing all eigenvalues through full SVD and then selecting the largest, TikUDA uses the power method to directly approximate the largest eigenvalue. This reduces the computational cost compared to full SVD, making the approach more efficient while ensuring effective alignment between source and target domains.
Figure 4: Diagram of the Proposed TikUDA
4 Experimental Setup
4.1 Case Study
@@
We assess our proposed method using two publicly available real-world datasets from IoT air quality sensor networks [39, 40]. The networks are located in Monte Cucco, an urban area in the province of Piacenza (Italy), designated as R-N69, and in Osio Sotto in Bergamo (Italy), designated as R-N212. Both networks were installed near reference stations operated by governmental agencies using high-precision instrumentation, providing a reliable ground truth for evaluation. Each network is equipped with a low-cost O3 sensor, NO2 sensor, temperature sensor, and relative humidity sensor. A detailed summary of the datasets is provided in Table 1 [11]. These two networks are separated by a considerable distance, leading to a distribution shift in sensor measurements, as illustrated in Figure 5. We evaluate the two datasets in both configurations: first using one dataset as the source and the other as the target, then reversing their roles. In Figure 6, we illustrate the effect of distribution shift on model performance in our experimental setup. As depicted, when the model is trained and deployed at the same measurement location, the performance remains competitive due to the fact that the distribution of the features remains constant. However, when the model is deployed at a different measurement location, a significant drop in performance is observed. This drop is attributed to the distribution shift between the source and target locations. We imitate a practical scenario where reliable ground truth measurements are not available at the target location. The task involves training a model on one labeled dataset (source) to predict reliable O3 and NO2 values at the target location.
Table 1: Summary of the Air Quality Datasets Used in the Study,
@@
R-69 O3, NO2, O3 (low-cost), NO2 (low-cost), Temperature, Relative Humidity OX-B431/NO2-B43F DHT1-Grove Monte Cucco 1 per hour 21/06/2018 - 25/09/2018
R-212 O3, NO2, O3 (low-cost), NO2 (low-cost), Temperature, Relative Humidity OX-B431/NO2-B43F DHT1-Grove Osio Sotto 1 per hour 27/06/2018 - 24/09/2018
@@
4.2 Data Preprocessing
@@ 4.3 Evaluation Metrics
$\mathsf{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$ (18)
$\mathsf{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$ (19)
In these equations, $n$ represents the total number of samples, $y_{i}$ is the actual observed value for the $i$-th estimation, and $\hat{y}_{i}$ is the predicted value for the $i$-th estimation. To provide a more comprehensive analysis, we evaluate the methods using both normalized and unnormalized labels. Normalized labels allow for results to be compared on a common scale, which is particularly useful when working with diverse datasets. Conversely, unnormalized labels retain the original scale of the data, providing insights into real-world applications.
4.4 Comparison Methods
@@
  • 6.
@@
4.5 Implementation Details
In this study, all compared methods share the same base architecture, except for AdaBN, which includes an additional batch normalization layer after the GAT layer to ensure a fair comparison. To determine the number of layers and the hidden dimension of the feature extractor, we split the source domain dataset into training and testing sets to perform model selection based on the labeled source data. This approach ensures that the feature extractor can effectively extract the necessary informative features for multisensor fusion and calibration tasks.
The feature extractor consists of four GRU layers and one GAT layer, with all layers (linear encoder, GRU, and GAT) set to a hidden dimension of 16. The remaining architecture includes a regressor, comprising one fully connected layer to predict the final output value. The model is updated using the Adam optimizer [44] with a learning rate of $3\times 10^{-4}$ and batch sizes of 64, over 150 epochs. For each training step, one batch of source and target domain data is processed.
Due to the absence of a labeled validation dataset, the hyperparameters and training specifications mentioned above were primarily chosen based on prior work [45]. In our approach, the hyperparameters $\gamma_{\text{angle}}$ and $\gamma_{\text{scale}}$, as defined in equation (17), were set to $\gamma_{\text{angle}}=10^{-2}\lambda$ and $\gamma_{\text{scale}}=10^{-3}\lambda$, respectively for ‘R-212 to R-69 (O3)’ scenario. Here, $\lambda=\frac{2}{1+\exp(-10p)}-1$ and $p$ represents the progress of training iterations, allowing $\lambda$ to range from 0 to 1. Consequently, $\gamma_{\text{angle}}$ and $\gamma_{\text{scale}}$ vary from 0 to $10^{-2}$ and $10^{-3}$, respectively. This follows a similar approach to [46, 47].
5 Results and Discussion
Table 2: RMSE Results for Four Domain Adaptation Scenarios
@@ 5 Results and Discussion
Table 3: MAE Results for Four Domain Adaptation Scenarios
@@
In this section, we assess the effectiveness of all compared methods based on their performance using RMSE and MAE metrics, as shown in Tables 2 and 3, respectively. The source-only model reveals that the prediction error for the O3 is more than four times larger than for the NO2 task in both transfer scenarios, indicating a significant domain shift in O3 estimation. Specifically, for the O3 task, our proposed method significantly improves RMSE from 59.40 (source-only model) to 20.35 (TikUDA) for the transfer task ‘R-212 to R-69’, and from 48.44 (source-only model) to 21.87 (TikUDA) for the transfer task ‘R-69 to R-212’. Overall, TikUDA demonstrates the best average performance compared to other methods across all adaptation and transfer scenarios, achieving an average RMSE of 0.105 and an average MAE of 0.078, outperforming all other methods.
DARE-GRAM shows the second-best performance with an average RMSE of 0.107 and an average MAE of 0.082, indicating that the proposed haversine distance function effectively helps align the source and target domains. Additionally, the proposed approximation of the Gram matrix using TikUDA did not negatively impact performance. Other methods such as MMD, CORAL, and ERM-NU also showed improvements in RMSE and MAE compared to the baseline, as indicated in Tables 2 and 3. However, they did not achieve the same level of performance, as they are not primarily designed for regression tasks like DARE-GRAM and TikUDA.
Additionally, we evaluate the impact of our approach on the embeddings generated during the alignment process for the adaptation scenario of R-212 $\rightarrow$ R-69 (O3). We applied principal component analysis (PCA) to the embedded features and plotted the first two principal components to create two-dimensional visualizations, as depicted in Figure 7. The PCA plot highlights a significant discrepancy between the source and target data distributions when using the source-only model. In contrast, our proposed method achieves a better alignment of data points from both domains, demonstrating its effectiveness in mitigating distribution shifts in the feature space.
(a) Baseline
(b) TikUDA
Figure 7: Visualization of the impact of the feature alignment with the proposed method on the (R-212 $\rightarrow$ R-69 (O3)) task
Furthermore, we compare the time complexity of our proposed algorithm with DARE-GRAM, as these two algorithms performed best in our experiments and share similar theoretical underpinnings, warranting a focused analysis. In other words, since these two methods are subspace-based approaches specifically designed for regression tasks, it is appropriate to present their running times side by side for a direct comparison. Both DARE-GRAM and TikUDA exhibit similar asymptotic computational complexities, approximately $\mathcal{O}(p^{3})$, but differ significantly in their practical execution. DARE-GRAM relies on full SVD and the computation of the Moore-Penrose pseudo-inverse, which leads to higher computational overhead. Specifically, calculating the Gram matrix for both methods requires $\mathcal{O}(bp^{2})$, where $b$ is the batch size and $p$ is the feature dimension. The SVD step further introduces an $\mathcal{O}(p^{3})$ complexity, and computing the pseudo-inverse also contributes an additional $\mathcal{O}(p^{3})$, making DARE-GRAM computationally expensive. In contrast, TikUDA utilizes power iteration to approximate the largest singular value, avoiding the need for full SVD. Each iteration of the power method takes $\mathcal{O}(p^{2})$. Moreover, TikUDA leverages Cholesky decomposition, which has a complexity of $\mathcal{O}(p^{3}/3)$, for matrix inversion, offering a more efficient alternative to SVD-based inversion. While both methods have similar theoretical upper bounds, the use of power iteration and Cholesky decomposition in TikUDA significantly reduce the constant factor in the complexity, leading to faster performance in practice compared to DARE-GRAM. Consequently, TikUDA is expected to perform more efficiently, especially for large-scale data, where full SVD can be prohibitively slow.
Figure 8: Comparison of running times on training iteration for TikUDA and DARE-GRAM across varying hidden dimensions ($128$ to $8192$).
5.1 Ablation Studies
In this subsection, we explore the effects of aligning angles and scaling eigenvalues on model performance through ablation studies. We select the tasks R-212 $\rightarrow$ R-69 (O3) and R-69 $\rightarrow$ R-212 (O3) for this analysis. As illustrated in Table 4, aligning both angles and scales in the proposed method results in performance enhancements compared to the source-only approach. Notably, scaling alignment significantly reduces the RMSE compared to the source-only baseline by ensuring the range of predicted values aligns across different datasets. Meanwhile, aligning angles alone yields a smaller improvement in performance.
Table 4: RMSE ($\downarrow$) Values for Ablation Studies
Method R-212 $\rightarrow$ R-69 (O3) R-69 $\rightarrow$ R-212 (O3)
Source-Only 0.255 0.217
$\gamma_{\text{angle}}=0$ 0.096 0.106
$\gamma_{\text{scale}}=0$ 0.202 0.159
Both (TikUDA) 0.087 0.099
The Tikhonov matrix introduces the hyperparameter $\alpha$. If $\alpha$ is too low, the matrix may not be fully ranked, leading to numerical instability. Conversely, a large $\alpha$ can cause the identity matrix subspace to become more prominent than that of the Gram matrix, which is undesirable for alignment. Our analysis of the R-212 $\rightarrow$ R-69 (O3) case, illustrated in Figure 9, shows that our method is relatively insensitive to variations in $\alpha$. This indicates that the specific choice of $\alpha$ is not critical to the performance of our proposed method.
Figure 9: Impact of Different Values of $\alpha$ on RMSE
6 Conclusion
This paper investigates the challenge of unsupervised domain adaptation regression within spatial-temporal graph neural networks, specifically focusing on the fusion and calibration of multisensor data in IoT air quality sensor networks. Our approach addresses domain adaptation for regression tasks by integrating the Tikhonov-regularized closed-form least square solution. Additionally, we introduce a novel method for scaling and aligning the Tikhonov matrix, which reduces complexity and improves accuracy. Experimental results demonstrate substantial improvements in the model’s generalization ability across four different real-world scenarios. The findings of this paper open up several potential directions for future research. Firstly, TikUDA is a model-agnostic method, indicating its applicability to various datasets and tasks across different fields, including computer vision, and multimodal learning. Moreover, future work can focus on exploring alternative regularization techniques, beyond ridge regression, that offer closed-form solutions for inverse problems. Implementing these techniques in unsupervised domain adaptation tasks holds promise for further advancements in the field.
Appendix A Details on Power Method
$\mathbf{v_{k}}=\frac{\mathbf{Av_{k-1}}}{\|\mathbf{Av_{k-1}}\|},$ (20)
[…]
Appendix B Generalization to Other Applications
Electroencephalography (EEG) Dataset: EEG signals are widely used in neuroscience and biomedical research due to their ability to capture the brain’s electrical activity in a non-invasive manner. The primary motivation for selecting an EEG dataset for this study lies in its inherent complexity, with significant variations in brain activity both within and across subjects. This complexity creates a challenging yet realistic setting for evaluating domain adaptation techniques. Due to the natural inter-subject variability in EEG signals, they serve as an ideal benchmark for evaluating our proposed TikUDA method. Effective knowledge transfer across domains, specifically between different subjects – is essential in applications such as brain-computer interfaces (BCIs) and neurological disorder diagnostics, where models must generalize across individuals with unique brainwave patterns.
The EEG recordings were obtained using a Neurocom 23-channel EEG system (Ukraine, XAI-MEDICA) configured in a monopolar setup [48]. Silver/silver chloride electrodes were positioned on the scalp following the International 10/20 system, with all electrodes referenced to interconnected ear reference electrodes. The data were processed using a high-pass filter with a 30 Hz cutoff and a 50 Hz power line notch filter. Artifact-free EEG segments, each lasting 60 seconds, were selected for analysis. During preprocessing, Independent Component Analysis (ICA) was employed to remove artifacts caused by eye movements, muscle activity, and cardiac pulsations. Participants engaged in a cognitive task involving the serial subtraction of two numbers, where each trial began with the oral presentation of a 4-digit minuend and a 2-digit subtrahend. The channels selected for analysis included Fp1, Fp2, F3, F4, F7, F8, T3, T4, C3, C4, T5, T6, P3, P4, O1, O2, Fz, Cz, and Pz as illustrated in Fig. 10
For the regression task, we focused on reconstructing the missing FP1 channel using the remaining active channels. In the domain adaptation task, we selected two subjects: Subject 00, a 21-year-old female, and Subject 33, a 17-year-old male. We assume that FP1 measurements are available for Subject 33, but for Subject 00, due to technical issues with the EEG cap, the FP1 channel was not recorded properly. This simulates a real-world scenario where sensor (electrode) malfunction or noise leads to missing data, and the goal is to leverage the data from other channels and the source subject (Subject 33) to accurately reconstruct the missing measurements for the target subject (Subject 00). This setting allows us to test the robustness of our method in handling missing data and domain shifts simultaneously.
Figure 10: EEG setup with 19 channels.
@@ Appendix B Generalization to Other Applications
We compared the same methods used in the air quality task and the STGNN architecture remained unchanged, allowing for a direct comparison of the approaches. The results in Table 5 highlight that in the domain adaptation scenario (S33 $\rightarrow$ S00) for EEG signals, while several methods perform well. However, TikUDA is consistently the best performing method. Specifically, while MMD and ERM-NU achieve competitive performance – with ERM-NU achieving an RMSE of 0.180 and an MAE of 0.132 – TikUDA outperforms these methods, achieving the lowest RMSE of 0.178 and the lowest MAE of 0.130. These findings demonstrate the effectiveness and robustness of TiKUDA in enhancing prediction accuracy for EEG-based domain adaptation tasks. Moreover, they confirm TiKUDA’s broad applicability to other UDA regression tasks across various domains.
Table 5: Results for Domain Adaptation Scenario (S33 $\rightarrow$ S00) on EEG Signals

### v2

@@ III-A TikUDA
[…]
@@
III-B Spatial-Temporal Feature Extractor
@@
Figure 4: Overview of the proposed approach for domain-adaptive regression on spatial-temporal graph data. The model extracts temporal features using GRU and spatial dependencies via GAT, followed by a regression head for target prediction. The TikUDA module enables unsupervised domain adaptation by aligning the source and target distributions.
Temporal Module of Feature Extractor : Temporal dependencies, also known as temporal patterns, can be effectively represented in both the time domain and the frequency domain [36]. To address these dependencies, we incorporate temporal modules before spatial modules, facilitating the modeling of complex spatial-temporal patterns. Our approach adopts recurrence-based methodologies, specifically leveraging the Gated Recurrent Unit (GRU), to capture and interpret the underlying temporal patterns inherent in the time series data.
@@
where $\mathbf{a}$ represents a trainable attention vector, $\mathbf{W}$ is a trainable matrix, and $||$ denotes the concatenation operation. For brevity, the $k$-th GAT layer can be represented compactly as:
@@
In terms of complexity, the GAT component scales as $\mathcal{O}(|\mathcal{V}|\cdot d^{2}+|\mathcal{E}|\cdot d)$ per attention head, where $|\mathcal{V}|$ is the number of nodes, $|\mathcal{E}|$ is the number of edges, and $d$ is the feature dimension. This accounts for both the linear transformations of node features ($\mathcal{O}(|\mathcal{V}|\cdot d^{2})$) and the attention coefficient computations over edges ($\mathcal{O}(|\mathcal{E}|\cdot d)$). For temporal modeling, GRU maintains a complexity of $\mathcal{O}(T\cdot N\cdot d^{2})$, where $T$ is the sequence length, $N$ is the number of nodes, and $d$ is the hidden size, due to matrix operations across all nodes and time steps.
III-C Overall Framework and Summary
@@
Next, the data undergoes a temporal processing step using GRU to identify and extract temporal patterns. This is followed by a spatial processing step employing a GAT. If a graph structure is available (e.g., EEG signal), the GAT operates on it directly; otherwise (e.g., air quality), we construct a fully connected graph to capture spatial dependencies. The extracted features are subsequently inputted into a regressor to estimate the target values.
In UDA procedure, TikUDA is designed to align the inverse of the Gram matrix of features, which is perturbed by a scaled identity matrix referred to as the Tikhonov matrix. This perturbation ensures that the Gram matrix becomes full rank and invertible. A key aspect of the method is that the Tikhonov matrix is both symmetric and positive definite, which allows for the use of Cholesky decomposition—a more computationally efficient approach to matrix inversion compared to SVD, particularly for large matrices. Cholesky decomposition, specifically suited for symmetric and positive definite matrices, enhances numerical stability while maintaining the structural properties of the Tikhonov matrix. Additionally, TikUDA introduces novel similarity metrics (Haversine similarity) as regularization terms to improve subspaces’ alignment between the source and target feature domains. Moreover, TikUDA focuses on scale alignment between source and target domains by aligning the largest principal eigenvalues of the inverse Tikhonov matrices, as the largest eigenvalue demonstrates the greatest influence on feature scaling. Instead of computing all eigenvalues through full SVD and then selecting the largest, TikUDA uses the power method to directly approximate the largest eigenvalue. This reduces the computational cost compared to full SVD, making the approach more efficient while ensuring effective alignment between source and target domains.
Finally, the training objective is to minimize both the prediction error on the source domain and the alignment of features between the source and target domains. The total loss function can be expressed as:
@@
The loss term $L_{\text{src}}$ represents the supervised loss (mean squared error) on the source domain, comparing the prediction with the ground truth. The terms $L_{\text{angle}}$ and $L_{\text{scale}}$ account for the angle and scale alignment with respect to the inverse Tikhonov matrix. The hyperparameters $\gamma_{\text{angle}}$ and $\gamma_{\text{scale}}$ control the influence of angle and scale alignment on the total loss.
IV Experimental Setup
IV-A Case Studies
@@
Air Quality Sensor Networks: We assess our proposed method using two publicly available real-world datasets from IoT air quality sensor networks [41, 42]. The networks are located in Monte Cucco, an urban area in the province of Piacenza (Italy), designated as R-N69, and in Osio Sotto in Bergamo (Italy), designated as R-N212. Both networks were installed near reference stations operated by governmental agencies using high-precision instrumentation, providing a reliable ground truth for evaluation. Each network is equipped with a low-cost O3 sensor, NO2 sensor, temperature sensor, and relative humidity sensor. A detailed summary of the datasets is provided in Table I [11]. These two networks are separated by a considerable distance, leading to a distribution shift in sensor measurements, as illustrated in Figure 5. We evaluate the two datasets in both configurations: first using one dataset as the source and the other as the target, then reversing their roles. In Figure 6, we illustrate the effect of distribution shift on model performance in our experimental setup. As depicted, when the model is trained and deployed at the same measurement location, the performance remains competitive due to the fact that the distribution of the features remains constant. However, when the model is deployed at a different measurement location, a significant drop in performance is observed. This drop is attributed to the distribution shift between the source and target locations. We imitate a practical scenario where reliable ground truth measurements are not available at the target location. The task involves training a model on one labeled dataset (source) to predict reliable O3 and NO2 values at the target location.
Electroencephalography (EEG) Case Study: EEG signals are widely used in neuroscience and biomedical research for their ability to non-invasively capture the brain’s electrical activity. We selected an EEG dataset for this study due to its inherent complexity and variability, both within and across individuals. This natural inter-subject variability presents a challenging yet realistic testbed for evaluating domain adaptation methods. Effective cross-subject knowledge transfer is essential in applications such as brain-computer interfaces and neurological disorder diagnostics, where models must generalize across individuals with distinct brainwave patterns – making EEG an ideal benchmark for our proposed TikUDA method.
The EEG data, obtained from [43], were recorded using a Neurocom 23-channel EEG system ( XAI-MEDICA, Ukraine) in a monopolar configuration. Electrodes were positioned according to the international 10/20 system and referenced to interconnected ear electrodes. Data preprocessing included a 30 Hz high-pass filter and a 50 Hz notch filter to remove power-line noise, as defined in previous studies [43]. Independent Component Analysis (ICA) was applied to remove artifacts caused by eye movements, muscle activity, and cardiac pulsations. Only artifact-free segments of 60 seconds were used for analysis. During each trial, participants engaged in a cognitive task involving the serial subtraction of a two-digit number from a four-digit number, presented orally at the start. For the regression task, we focused on reconstructing the FP1 channel using the remaining active channels: Fp1, Fp2, F3, F4, F7, F8, T3, T4, C3, C4, T5, T6, P3, P4, O1, O2, Fz, Cz, and Pz, as illustrated in Fig. 7 In the domain adaptation setup, we selected two subjects: Subject 33 (a 17-year-old male) as the source and Subject 00 (a 21-year-old female) as the target. Due to a malfunctioning EEG cap, the FP1 channel was not properly recorded for for Subject 00, simulating a real-world scenario of sensor failure or noise. The task, therefore, involves leveraging data from other channels and the source subject to reconstruct missing measurements for the target subject – testing the model’s robustness to both missing data and inter-subject domain shifts.
TABLE I: Summary of the Air Quality Datasets Used in the Study,
@@
R-69
O3, NO2, O3 (low-cost), NO2 (low-cost),
Temperature, Relative Humidity
OX-B431/NO2-B43F
DHT1-Grove
Monte Cucco
1 per hour
21/06/2018 - 25/09/2018
R-212
O3, NO2, O3 (low-cost), NO2 (low-cost),
Temperature, Relative Humidity
OX-B431/NO2-B43F
DHT1-Grove
Osio Sotto
1 per hour
27/06/2018 - 24/09/2018
@@
Figure 7: EEG setup with 19 channels.
IV-B Data Preprocessing
@@ IV-C Evaluation Metrics
To provide a more comprehensive analysis, we evaluate the methods using both normalized and unnormalized labels. Normalized labels allow for results to be compared on a common scale, which is particularly useful when working with diverse datasets. Conversely, unnormalized labels retain the original scale of the data, providing insights into real-world applications.
IV-D Comparison Methods
@@
  • Adaptive Graph Convolutional Network (AdaGCN): AdaGCN [47] leverages adversarial domain adaptation and graph convolution to transfer label information from labeled source data to unlabeled target data. It uses an adversarial domain adaptation component that mitigates the distribution shift between the source and target domains.
@@
IV-E Implementation Details
In this study, all compared methods share the same base architecture, except for AdaBN, which includes an additional batch normalization layer after the GAT layer to ensure a fair comparison. To determine the number of layers and the hidden dimension of the feature extractor, we split the source domain dataset into training and testing sets to perform model selection based on the labeled source data. This approach ensures that the feature extractor can effectively extract the necessary informative features for multisensor fusion and calibration tasks. The feature extractor consists of four GRU layers and one GAT layer, with all layers (linear encoder, GRU, and GAT) set to a hidden dimension of 16. The remaining architecture includes a regressor, comprising one fully connected layer to predict the final output value. The model is updated using the Adam optimizer [48] with a learning rate of $3\times 10^{-4}$ and batch sizes of 64, over 150 epochs. For each training step, one batch of source and target domain data is processed.
In UDA, the absence of labeled data for the target domain makes it impossible to utilize a validation dataset for hyperparameter tuning. Consequently, an alternative approach must be adopted. We employed a heuristic method proposed in [49], which addresses this challenge. Our observations indicated that if the trade-off alignment regularization is set too high for all methods, it hinders the model’s ability to learn effectively. Conversely, if it is too low, there is insufficient adaptation to the target domain. To tune the hyperparameters, we conducted a search over specific values and tracked the normalized summation of the source loss and the trade-off alignment loss. This process was applied to all methods to ensure a fair comparison. In our approach, the hyperparameters $\gamma_{\text{angle}}$ and $\gamma_{\text{scale}}$, as defined in equation (13), were set to $\gamma_{\text{angle}}=10^{-2}\lambda$ and $\gamma_{\text{scale}}=10^{-3}\lambda$, respectively for the first scenario (air quality case study). Here, $\lambda=\frac{2}{1+\exp(-10p)}-1$ and $p$ represents the progress of training iterations, allowing $\lambda$ to range from 0 to 1. Consequently, $\gamma_{\text{angle}}$ and $\gamma_{\text{scale}}$ vary from 0 to $10^{-2}$ and $10^{-3}$, respectively. This follows a similar approach to [50, 51]. For EEG data, hyperparameters were set to $\gamma_{\text{angle}}=5\times 10^{-2}\lambda$ and $\gamma_{\text{scale}}=2.5\times 10^{-3}\lambda$.
V Results and Discussion
TABLE II: RMSE Results for Four Domain Adaptation Scenarios
@@ V Results and Discussion
TABLE III: MAE Results for Four Domain Adaptation Scenarios
@@
In this section, first, we assess the effectiveness of all compared methods based on their performance using RMSE and MAE metrics for air quality dataset, as shown in Tables II and III, respectively. The source-only model reveals that the prediction error for the O3 is more than four times larger than for the NO2 task in both transfer scenarios, indicating a significant domain shift in O3 estimation. Specifically, for the O3 task, our proposed method significantly improves RMSE from 59.40 (source-only model) to 20.35 (TikUDA) for the transfer task ‘R-212 to R-69’, and from 48.44 (source-only model) to 21.87 (TikUDA) for the transfer task ‘R-69 to R-212’. Overall, TikUDA demonstrates the best average performance compared to other methods across all adaptation and transfer scenarios, achieving an average RMSE of 0.105 and an average MAE of 0.078, outperforming all other methods. DARE-GRAM shows the second-best performance with an average RMSE of 0.107 and an average MAE of 0.082, indicating that the proposed haversine distance function effectively helps align the source and target domains. Other methods such as MMD, CORAL, ERM-NU, and AdaGCN also showed improvements in RMSE and MAE compared to the baseline, as indicated in Tables II and III. However, they did not achieve the same level of performance, as they are not primarily designed for regression tasks like DARE-GRAM and TikUDA.
@@
For the EEG dataset, we employed the same set of domain adaptation methods as in the air quality task, using an identical STGNN architecture. This consistency ensures a fair and direct comparison across domains. The results in Table IV show that while several methods perform well in the domain adaptation scenario (S33 $\rightarrow$ S00), TikUDA consistently achieves the best performance. Specifically, although AdaGCN and ERM-NU achieve competitive results, TikUDA surpasses both, achieving the lowest RMSE of 0.178 and the lowest MAE of 0.130. These results demonstrate TikUDA’s effectiveness and robustness in EEG-based domain adaptation, particularly in handling inter-subject variability and missing data. Furthermore, they underscore TiKUDA’s broad applicability to UDA regression tasks across diverse domains.
TABLE IV: Results for Domain Adaptation Scenario (S33 $\rightarrow$ S00) on EEG Signals